Two investments are made totaling $14,000. For a certain year these investments yield $1064 in simple interest. Part of the $14,000 is invested at 6% and the rest at 8%. How much is invested at 8%?
Answer by hamsanash1981@gmail.com(151)

Let the amount invested at 8% be x then the amount invested at 6% is 14000-x
then
SI at 6% = (14000-x)*6/100 = (84000-6x)/100
SI at 8% = 8x/100
8x/100 + (84000-6x)/100 = 1064
2x +84000 =106400
2x = 106400-84000
2x = 22400
x = 11200
The amount invested at 8% is 11200.